OpenStudy (anonymous):

2(n-3)+7=3(n-1) Solve for n. I multiplied the numbers in the parenthesis, but I am not sure what to do next. PLZ Help

jimthompson5910 (jim_thompson5910):

2(n-3)+7=3(n-1) 2n-2*3+7=3n-3*1 ... distribute 2n-6+7=3n-3 ... multiply 2n+1=3n-3 ... combine like terms. I'll let you finish up.

OpenStudy (anonymous):

is 2 or -2

OpenStudy (anonymous):

or is 4